Solve the following system of equations.


5x - y = 7
2 + x = 7y

Answer by Cromlix(4381)

5x - y = 7 ....1
x - 7y = -2 ....2
Multiply (2) by -5
5x - y = 7
-5x + 35y = 10
Add together
34y = 17
y = 17/34 or 1/2
Substitute y = 1/2 in (1)
5x - 1/2 = 7
5x = 7 + 1/2
5x = 15/2
x = 15/10 or 3/2
{3/2, 1/2}